OpenStudy (anonymous):

What is the point slope form of (60,58)?

OpenStudy (anonymous):

must be a bit more to the question, right?

OpenStudy (anonymous):

I know the formula is y - y1 = m(x - x1)

OpenStudy (anonymous):

sub in the x and the y into the formula

OpenStudy (anonymous):

do you have the m? the slope?

OpenStudy (anonymous):

I did and I got y - 58 = 1.06(x - 60) 1.06 is the slope

OpenStudy (anonymous):

You will need one more piece of information to put this into y-intercept form or standard form

OpenStudy (anonymous):

you only have one point. you need either one more point or a slope. there is no "point-slope" formula for one point

OpenStudy (anonymous):

1.06 is the slope

OpenStudy (anonymous):

whit what info you have given it would be y=1.06x-5.6

OpenStudy (anonymous):

Nulify, just give out the original question! We don't want to misguide you!